i need to know how to solve and understand this problem: 2√126/√14
Answer by benazir.sj@gmail.com(109)

2√126/√14
solution:
√126=√2*3*3*7
= 3√14
2√126/√14 =2*3√14/√14 (since√14 is common we can cancel )
=6